The perimeter of a quadrilateral is 38 centimeters. Find the length of each side.
x+8
x+4
x^2-2x
3x-2

Answer by ikleyn(52781)

.
Perimeter = 38 = (x+8) + (x+4) + (x^2-2x) + (3x-2)


38 = x^2 + 3x - 2

x^2 + 3x - 40 = 0

(x+8)*(x-5) = 0


The only positive root is x= 5.


The sides are  5+8= 13;  5+4 = 9;  5^2-2*5 = 15;  and  3*5-2 = 13.

Solved.